OpenStudy (anonymous):

Which one of the following sets of numbers represents the lengths of the sides of a right triangle?

OpenStudy (anonymous):

A. 6, 9, 10 B. 9, 40, 41 C. 5, 10, 15 D. 13, 35, 37

OpenStudy (anonymous):

Shouldn't a angle be 90?

OpenStudy (anonymous):

the one where \(a^2+b^2=c^2\)

OpenStudy (anonymous):

Isn't that the Pythagorean formula?

OpenStudy (anonymous):

Taking C for example 25+100=225

OpenStudy (anonymous):

So not C right @satellite73 ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Ok, I got this. Thanks! :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

hehe I got it! B!

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks again.